# Large Language Model Lateral Spear Phishing: A Comparative Study in Large-Scale Organizational Settings

### v1

@@ Abstract
The critical threat of phishing emails has been further exacerbated by the potential of LLMs to generate highly targeted, personalized, and automated spear phishing attacks. Two critical problems concerning LLM-facilitated phishing require further investigation: 1) Existing studies on lateral phishing lack specific examination of LLM integration for large-scale attacks targeting the entire organization, and 2) Current anti-phishing infrastructure, despite its extensive development, lacks the capability to prevent LLM-generated attacks, potentially impacting both employees and IT security incident management. However, the execution of such investigative studies necessitates a real-world environment, one that functions during regular business operations and mirrors the complexity of a large organizational infrastructure. This setting must also offer the flexibility required to facilitate a diverse array of experimental conditions, particularly the incorporation of phishing emails crafted by LLMs. This study is a pioneering exploration into the use of Large Language Models (LLMs) for the creation of targeted lateral phishing emails, targeting a large tier 1 university’s operation and workforce of approximately 9,000 individuals over an 11-month period. It also evaluates the capability of email filtering infrastructure to detect such LLM-generated phishing attempts, providing insights into their effectiveness and identifying potential areas for improvement. Based on our findings, we propose machine learning-based detection techniques for such emails to detect LLM-generated phishing emails that were missed by the existing infrastructure, with an F1-score of 98.96. Our findings also highlight the urgent need for integrating existing anti-phishing infrastructure with LLM-generated phishing email detection methods and point out the need for updated organizational policies towards mitigating LLM driven phishing threats.
1 Introduction
@@
The rapid advancement of internet technologies has significantly heightened security concerns, especially in the realm of phishing, a critical form of social engineering that deceitfully extracts sensitive information [1], [2], [3], [4]. Attackers masquerade as legitimate entities to gain unauthorized access to accounts and acquire confidential data [5]. The education sector, in particular, has been heavily targeted, experiencing a 576% increase in phishing attacks from 2021 to 2022 [6]. Lateral spear phishing, which exploits both technical and human vulnerabilities by crafting personalized communications, has become a sophisticated and deceptive attack method [7]. This landscape is further complicated by the emergence of Large Language Models (LLMs) like OpenAI’s ChatGPT, which have shown remarkable capabilities in generating human-like text [8], [9], [10]. These developments transform traditional phishing tactics, making them more nuanced and context-aware. Cisco Systems Inc. has highlighted the increased difficulty in detecting phishing attempts due to AI, urging the adoption of new defenses [11]. Additionally, the use of AI in phishing has been noted to enhance the frequency and sophistication of social engineering attacks [12]. In spearphishing, the advanced capabilities of LLMs to tailor messages to individual targets could significantly increase the threats’ sophistication and effectiveness [13].
@@
This escalating threat landscape, particularly with the involvement of LLMs in phishing, necessitates innovative studies to understand this threat and the exploration of novel cybersecurity defenses. There are two main problems regarding this new threat. First, existing studies on lateral phishing attacks do not investigate the integration of LLMs for large-scale attacks targeting the entire organizations. Previous works that analyze phishing attacks at scale, provide valuable insight into the tactics and techniques as well as mitigation strategies towards phishing threats [14, 15, 16], but do not consider the role that LLMs could play in this threat. However, conducting such studies demands a real-world environment that is representative of a large-scale organizational structure. Second, existing anti-phishing infrastructure lack the capability to prevent LLM-generated attacks, potentially impacting both employees and IT operations support and incident management. Traditional phishing filters rely on rule-based methods of identifying phishing by identifying known phishing sites and senders [17]. However, the test environment used to investigate the feasibility of using automatic defense methods against LLM-generated attacks must be flexible enough to incorporate both LLM-generated attacks and defenses involving lateral phishing red-teaming using LLMs, which presents a novel and highly effective approach in cybersecurity exercises, aimed at enhancing organizational defenses against sophisticated phishing threats.
Our investigation, situated within a large public university, delves into the specific cyber threats faced by the institution, with a particular focus on the use of large language models (LLMs) in generating lateral phishing content. In our study, we conducted a large scale study across roughly 9000 employees over the period of 11 months where we created a variety of different phishing email templates. The existing phishing simulation infrastructure of the university’s cyber operations team supports the logging of information of how users interact with simulated phishing emails, collecting information such as if they open the email, click the phishing link or input login credentials. Using this infrastructure, we investigate both human written and LLM generated phishing emails, examining how internal and external organizational information can be used to craft effective lateral spear phishing attacks by capturing the data on how users interact with these simulated phishing attacks.
This comprehensive study not only examines the efficacy of LLM-crafted phishing emails, which we found to be strikingly similar to traditional human-generated lateral spear phishing attacks, but also incorporates valuable insights from the university’s cyber operations team. In one of our critical experiments, it was observed that about 10% of email recipients at the university were compelled to input their login credentials when targeted by LLM-generated phishing emails. This underscores the persuasive effectiveness of these AI-crafted attacks. Our research further explores the factors that heighten vulnerability to these sophisticated attacks and pinpoints the elements within the phishing emails that impact recipients’ decision-making, especially in terms of whether to engage with or dismiss these emails. Notably, the use of information specific to the internal versus external dynamics of the organization proved to be a significant factor in the success rate of these LLM-driven lateral spear phishing attempts. Additionally, we present insights from the organization’s cyber operations team, who emphasize the need for updated phishing trainings and warnings to better address the nuances of LLM phishing threats. Their expert perspectives shed light on potential mitigation strategies and the necessity of evolving educational approaches to keep pace with these emerging cyber threats. To combat these sophisticated attacks, our study further showed advanced machine learning-based detection techniques designed specifically to identify LLM-generated phishing emails. These methods have achieved a high success rate, with an F1-score of 98.96, successfully identifying phishing attempts that had bypassed existing security protocols. The integration of these novel detection techniques into the university’s existing anti-phishing infrastructure is highlighted as a potential next step towards enhancing the organization’s resilience against these increasingly complex and evolving cyber threats.
@@
  • We undertake a thorough study to show the state of cyber security threats at a large public university. Alongside this we show how insights from the organization’s cyber operations team drive our investigation into LLM generated spear phishing attacks.
@@
  • We conduct extensive experiments spanning 11 months across roughly 9000 employees at a large public university to analyze how they interact with phishing emails. We further show how LLM generated lateral spear phishing emails utilizing internal information can is a significant security risk.
  • We develop and a large-scale detection framework for identifying emails generated by large language models, demonstrating rapid detection capabilities with minimal operational overhead.
2 Organization Security Incident Investigation: A Motivational Study
@@
To contextualize our focus on studying LLMs for lateral spear phishing, we initially delve into the background of the organization central to this research. Following this, we outline the cyber security threats it encounters, highlighting phishing as a predominant and financially damaging cyber threat. Subsequently, we detail a specific instance of lateral spear phishing that the organization experienced earlier in the year. Building on this, we draw upon insights from the organization’s cyber operations team, which inform the design of our phishing experiments and shape the research questions we aim to answer through this study.
2.1 Organization Background
The organization has a workforce of approximately 9,000 employees and contributes approximately \$2.5 billion to the local economy. Its customer segment, predominantly students, numbers just under 35,000. Spanning over 700 acres, the main campus encompasses four distinct campuses and features over sixty buildings. The IT organization delivers enterprise-level services within a multifaceted and federated environment, comprising both centralized and distributed IT operations, with the majority of services being managed by the central IT team. This department adheres to a well-established ITIL framework and primarily aligns with NIST security standards to meet research and educational regulatory requirements. The network, both wired and wireless, regularly supports over 70,000 connected devices. The cyber operations team operates using the MITRE framework, conducting monthly phishing simulation campaigns, weekly distributions of scam alerts for educational purposes, and collaborating with external groups for penetration testing exercises (blue team versus red team scenarios). The team is bolstered by senior-level students who participate as interns or work-study members, gaining practical experience in the field. All first-level IT support is centralized in a large unit, equipped to manage calls, incidents, and requests from all campus locations. The diagram in Figure 1 illustrates the deployment of lateral phishing exercises and how it takes place within the organization.
2.2 Lateral Phishing and Impact
@@
Figure 2: Diverse participation in red-teaming exercises across various departments and job functions within the organization.
To enhance the relevance of our study on lateral phishing emails, we also underscore a recent spear phishing incident within our organization. In this lateral spear phishing incident, seven accounts were impacted, resulting in a combined effort of more than 2,000 soft labor hours and an estimated cost of \$73,500 to the organization. In the first half of 2023, a lateral spear phishing operation set a precedent for our phishing exercises in October and November. This attack specifically targeted high-ranking individuals, employing a sense of urgency in the email content and using cunning methods to propagate the attack, including creating inbox rules so the attacks would hide in plain sight. The primary aim was to harvest credentials, with a focus on diverting financial transactions to an overseas account. This sophisticated method underscored the urgent need to improve our defenses against such attacks. It also provided a valuable opportunity to assess the effectiveness of our current training protocols. In response, our organization revised its monthly training modules and invested in an advanced phishing detection system that utilizes behavioral analytics to more efficiently detect various phishing schemes.
@@
2.3 Social Engineering Observations
@@
Social engineering tactics play a pivotal role in the effectiveness of phishing campaigns, as observed by our organization’s cyber operations team. These insights reveal how attackers manipulate human psychology, often through seemingly legitimate offers or urgent requests, to trick individuals into compromising their security. One significant trend observed in lateral phishing attacks within the organization involves targeting recent contacts, especially direct reports, when the compromised account belongs to a supervising employee, validating observations on previous research on lateral phishing attacks [18, 14]. This method adds an insidious layer to the attack, leveraging the existing trust within professional hierarchies. Another notable example observed by the cyber operations team is phishing emails disguised as job opportunities, primarily targeting students. These emails cunningly use external links or attached Word documents as bait. In a broader context, phishing emails regarding Duo MFA, credential renewals, or resets are crafted with urgency to deceive recipients across the organization. Another widespread tactic involves falsely alarming users about the de-provisioning of their email accounts if they fail to take immediate action. Faculty and staff are often the targets of these sophisticated phishing attempts. These include spoofed login pages from recognized domains (such as r20.rs6.net) and emails with forged usernames from external domains, employing spearphishing strategies. To enhance their deceptive appearance, many phishing emails include images linked to the organization or Microsoft, exploiting the trust associated with these entities. This integration of social engineering techniques into phishing campaigns underscores the need for continual vigilance and education in cyber security practices.
The cyber operations team also discussed the potential threat of LLMs toward generating highly sophisticated spear phishing emails. Traditional phishing filters, which predominantly rely on rule-based systems or databases of known phishing sites and senders [17, 19], are finding it increasingly difficult to detect these AI-generated emails. The reason is that these emails often bypass standard detection rules by avoiding typical phishing indicators such as whether the email is sent by an unusual sender, and they are not yet listed in existing databases of known threats. As a result, these emails, which can be highly personalized and contextually relevant, pose a significant threat as they are difficult to distinguish from genuine communications. To effectively counter this new wave of AI-generated spear phishing, they expressed a need for more advanced solutions. These include AI-driven filters that can understand context and intent, behavior-based analysis, and continuous learning systems that adapt to evolving threats. Additionally, human vigilance remains a critical component. Educating users about the sophistication of these attacks and fostering a culture of skepticism and caution is essential in mitigating the risk posed by these advanced phishing attempts.
2.4 Security Incident Conclusion
The phishing email templates created for our research are significantly influenced by the expert insights of our organization’s cyber operations team, particularly regarding the dynamics of supervisor-direct report relationships in phishing attacks. These insights have led to the development of templates that mimic this specific relationship, using elements of urgency and personalization to create a sense of legitimacy and prompt action. This approach is grounded in the team’s observations of past successful phishing campaigns, where the exploitation of the hierarchical trust in an organization proved to be a particularly effective strategy.
Moreover, the team’s analysis highlighted the potential vulnerability of our organization to phishing attacks that use publicly available information. This understanding inspired the creation of a novel template that does not rely on the supervisor-direct report dynamic. Instead, it utilizes information that is publicly accessible external information about the organization. It serves as a crucial indicator of how well our organization can withstand phishing attempts that are based on public knowledge, rather than personal relationships, which may become more effective and prevalent as LLMs are able to generate more convincing and contextually relevant content. As these models become increasingly sophisticated in synthesizing and utilizing publicly available information, they can create phishing messages that are highly tailored and believable, potentially bypassing traditional security awareness trained on recognizing more personalized, direct-report based lateral phishing attempts.
@@ 2.4 Security Incident Conclusion
  • Spearphishing emails often mimic communications from a supervisor to their direct reports, presenting a deceptive appearance of legitimacy.
@@
  • The crafting of spear phishing emails may involve the use of LLMs.
2.5 Research Questions
RQ1: How effective are lateral phishing attacks across a large educational organization? Understanding the effectiveness of lateral phishing attacks is critical to assessing organizational security vulnerabilities. This research question is novel as it investigates a large educational institution, contrasting with previous large-scale phishing research which predominantly focuses on commercial organizations. Our approach offers a unique perspective compared to the findings presented in previous prominent large scale phishing studies [14, 15]. The first part of this study involves multiple phishing email campaigns across 11 months where we report statistics on how users engage with the simulated phishing emails.
RQ2: Do LLM crafted lateral phishing emails have similar impact as human-generated lateral phishing emails? This question is essential to evaluate the evolving threats in cybersecurity, particularly the effectiveness of AI in mimicking human phishing tactics. In the final month of this phishing study, a similar phishing exercise is conducted using LLM-generated emails, allowing for a comparison of the impact between AI-generated and human-generated phishing attempts.
RQ3: How does the effectiveness of LLM-generated spear phishing emails differ when using internal versus external organizational information? This question is crucial as it explores the potential threat level posed by phishing attacks crafted with varying degrees of insider knowledge. Specifically, we investigate the efficacy of spear phishing attacks that utilize information about an organization’s structure and internal details compared to those based solely on publicly available information. The study involves an analysis where an LLM is used to create phishing emails. One set of emails is generated using data accessible from the organization’s public website, while another set leverages more detailed internal information.
RQ4: How effective are phishing training and warnings in lateral phishing across an organization? Evaluating the effectiveness of existing phishing countermeasures is crucial for improving organizational security protocols [20]. The study involves a targeted analysis of individuals who received phishing training, especially those who previously entered credentials in a phishing attack, to measure the training’s impact on reducing susceptibility to phishing.
RQ5: Can LLM detectors be used to identify LLM-generated phishing emails? Investigating the effectiveness of LLM detectors in identifying AI-generated emails is vital for analyzing how LLM generated phishing emails may be combated. The research includes an analysis where human generated emails are regenerated using an LLM and tested against a handful of machine text detectors. Both existing machine text detectors and a system using LLM embeddings are tested for their effectiveness in distinguishing machine-generated emails.
3 Threat Model

### v2

@@ Abstract
The emergence of Large Language Models (LLMs) has heightened the threat of phishing emails by enabling the generation of highly targeted, personalized, and automated attacks. Traditionally, many phishing emails have been characterized by typos, errors, and poor language. These errors can be mitigated by LLMs, potentially lowering the barrier for attackers. Despite this, there is a lack of large-scale studies comparing the effectiveness of LLM-generated lateral phishing emails to those crafted by humans. Current literature does not adequately address the comparative effectiveness of LLM and human-generated lateral phishing emails in a real-world, large-scale organizational setting, especially considering the potential for LLMs to generate more convincing and error-free phishing content. To address this gap, we conducted a pioneering study within a large university, targeting its workforce of approximately 9,000 individuals including faculty, staff, administrators, and student workers. Our results indicate that LLM-generated lateral phishing emails are as effective as those written by communications professionals, emphasizing the critical threat posed by LLMs in leading phishing campaigns. We break down the results of the overall phishing experiment, comparing vulnerability between departments and job roles. Furthermore, to gather qualitative data, we administered a detailed questionnaire, revealing insights into the reasons and motivations behind vulnerable employee’s actions. This study contributes to the understanding of cyber security threats in educational institutions and provides a comprehensive comparison of LLM and human-generated phishing emails’ effectiveness, considering the potential for LLMs to generate more convincing content. The findings highlight the need for enhanced user education and system defenses to mitigate the growing threat of AI-powered phishing attacks.
Index Terms:
Artificial intelligence, cybersecurity, disinformation, generative AI, large language models, phishing, text generation
[…]
=-21pt
I Introduction
@@
The rapid advancement of internet technologies has significantly heightened security concerns, especially in the realm of phishing, a critical form of social engineering that deceitfully extracts sensitive information [1], [2], [3], [4]. Attackers masquerade as legitimate entities to gain unauthorized access to accounts and acquire confidential data [5]. The education sector in particular has been heavily targeted [6], and in 2024 continues to be the most heavily targeted sector [7]. Lateral phishing, which exploits both technical and human vulnerabilities by crafting personalized communications, has become a sophisticated and deceptive attack method [8]. While knowledgeable users may not be easily deceived by less sophisticated attackers [9], whose poor grammar and obvious inconsistencies give them away, the advent of Large Language Models (LLMs) like OpenAI’s ChatGPT has complicated the landscape. By producing remarkably human-like text [10], [11], [12], these LLMs can empower even less advanced attackers, significantly broadening their capabilities to mount more sophisticated and convincing phishing attacks. Additionally, the use of AI in phishing has been noted to enhance the frequency and sophistication of social engineering attacks [13].
@@
This escalating threat landscape, particularly with the involvement of LLMs in phishing, necessitates innovative studies to understand this threat. There are three main problems regarding the existing literature with respect to this new threat. First, existing studies on lateral phishing attacks do not investigate the integration of LLMs for large-scale attacks targeting entire organizations. Previous works that analyze phishing attacks at scale, provide valuable insight into the tactics and techniques as well as mitigation strategies towards phishing threats [14, 15, 16], but do not consider the role that LLMs could play in this threat. However, conducting such studies demands a real-world environment that is representative of a large-scale organizational structure. Second, while some recent work has been done towards the automated detection of LLM generated phishing emails [17], there has yet to be work to quantify users’ vulnerability to such emails on a large scale. Third, there’s a notable gap in understanding the reasons why individuals might interact with lateral phishing emails, as there is a lack of studies directly asking those who fell for such attacks about their motivations and thought processes.
In this research, we aim to answer the following research questions: 1.) How does the effectiveness of lateral phishing emails generated by Large Language Models compare to that of those crafted by humans within an educational organization? 2.) Which job roles and department affiliations exhibit the highest risk of falling victim to phishing emails, as evidenced by elevated click-through rates, and what patterns can be discerned to identify these high-risk groups? 3.) What are the underlying reasons for individuals to engage with lateral phishing emails, as revealed through questionnaires with those who have clicked on them? Our investigation, situated within a large public university, first delves into the specific cyber threats faced by the institution. Investigating the history of security incidents within the organization and through discussions with the cyber operations team, we find that phishing campaigns are the most common security incident, with a 15% increase in just the past year and that LLMs may be to blame. Informed by these observations, we conducted a large scale study across roughly 9000 employees where we tested a handful of different phishing email templates. The existing phishing simulation infrastructure of the university’s cyber operations team supports the logging of information of how users interact with simulated phishing emails, collecting information such as if they open the email, click the phishing link or input login credentials. Furthermore, we conduct a questionnaire on those who input login credentials in the phishing experiment to understand why users interact with the phishing emails, in order to understand some of the key decision-making factors that lead individuals to fall prey to these sophisticated cyber threats, thereby informing more effective user education and system defenses. Our study shows that LLM-generated lateral phishing emails are as effective as, and in some cases even more effective than, those crafted by human communications professionals. Notably, in one of the most effective lateral phishing scenarios, where the sender is a compromised account of a supervisor, LLM-generated emails performed comparably to human-crafted ones, exploiting trust relationships and power dynamics within the organization. Furthermore, LLM-generated emails significantly outperformed human-crafted emails in the timely phishing scenario, capitalizing on current events and institutional context. Our analysis also identified that student roles exhibited the highest vulnerability to phishing attempts, with the highest percentages of both links clicked and data entered compared to other job roles. Additionally, insights gathered from a questionnaire administered to employees who entered data in the simulated phishing emails revealed that the sender’s identity, perceived relevance of the email, and the sense of urgency were key factors influencing recipients’ decision to interact with the phishing content.
@@
  • We undertake a thorough study to show the state of cyber security threats at a large public university. Alongside this we show how insights from the organization’s cyber operations team drive our investigation into LLM generated phishing attacks.
@@
  • We conduct experiments across roughly 9000 employees at a large public university to compare how they interact with LLM written phishing emails versus human written phishing emails.
  • We administered a detailed questionnaire to 34 employees who entered data into the simulated phishing emails, aiming to understand the reasons behind their actions. This allowed us to gather insights, which can inform future strategies for phishing mitigation within the organization.
II Related Work
Recently, concerns around LLMs being used for adversarial purposes, particularly in the context of phishing attacks, have risen significantly. A recent report by Google Cloud underscores that LLMs can be used to create highly legitimate-seeming content, including voice and video, making it more challenging to identify misspellings, grammar errors, and cultural inaccuracies in phishing emails and messages [29]. There has also been an observed increase in the volume of phishing attackcs since the launch of ChatGPT, however the attack tactics and types of attacks have remained the same [30]. LLMs can also effectively translate and refine content, further complicating the detection of phishing attempts based on language use. The report points out that with generative AI, attackers can scale their operations and target a wide range of individuals with personalized and convincing emails, using data like names, organizations, job titles, and even health information. The work of Hazell et al. studied the effectiveness of LLMs toward creating realistic and cost-effective spear phishing emails to over 600 British Members of Parliament [31]. Their work also notes the possibility of bypassing safeguards in LLMs through basic prompt engineering, and emphasize the need for robust governance measures to prevent misuse. Another recent study [32] involved involved two groups of participants, exposed to either human-crafted or GPT-3 crafted phishing emails. Participants were asked to determine the authenticity of the emails, with feedback provided in the second round. The results indicated that human-crafted emails were more effective in deceiving people than those created by GPT-3, even after participants received training across various cognitive biases. Interestingly, participants felt more confident identifying phishing emails when they were human-crafted compared to those generated by GPT-3. Some recent work also explored how LLMs can be used to rewrite phishing emails to help bypass traditional email phishing filters [33]. Beyond phishing emails, there has also been work that utilizes LLMs to generate the source code of phishing websites [34]. As a response to these growing concerns, there has been an increased focus on detecting LLM-generated content [35, 36, 37]. Additionally, researchers are exploring the use of LLMs themselves to help detect phishing attempts [38].
III Organization Security Incident Investigation: A Motivational Study
@@
To contextualize our focus on studying LLMs for lateral phishing, we outline the cyber security threats the organization encounters based on official incident data provided by the organization’s Cyber Operations Team, highlighting phishing as a predominant and financially damaging cyber threat. Subsequently, we detail a specific instance of lateral phishing that the organization experienced earlier in the year, documented in the team’s security records. Building on this empirical evidence, we draw upon quantitative insights and threat intelligence from the organization’s cyber operations team, which inform the design of our phishing experiments and shape the research questions we aim to answer through this study.
III-A Lateral Phishing and Impact
@@
To enhance the relevance of our study on lateral phishing emails, we also underscore a recent phishing incident within our organization. In this lateral phishing incident, seven accounts were impacted, resulting in a combined effort of more than 2,000 soft labor hours and an estimated cost of \$73,500 to the organization. This attack specifically targeted high-ranking individuals, employing a sense of urgency in the email content and using cunning methods to propagate the attack, including creating inbox rules so the attacks would hide in plain sight. The primary aim was to harvest credentials, with a focus on diverting financial transactions to an overseas account. This sophisticated method underscored the urgent need to improve defenses against such attacks. It also provided a valuable opportunity to assess the effectiveness of our current training protocols. In response, the organization revised its monthly training modules and invested in an advanced phishing detection system that utilizes behavioral analytics to more efficiently detect various phishing schemes.
@@
III-B Social Engineering Observations
@@
Social engineering tactics play a pivotal role in the effectiveness of phishing campaigns, as observed by our organization’s cyber operations team. These insights reveal how attackers manipulate human psychology, often through seemingly legitimate offers or urgent requests, to trick individuals into compromising their security. One significant trend observed in lateral phishing attacks within the organization involves targeting recent contacts, especially direct reports, when the compromised account belongs to a supervising employee, validating observations on previous research on lateral phishing attacks [39, 14]. This method adds an extra layer to the attack, leveraging the existing trust within professional hierarchies. Another observation in existing phishing attacks faced by the organization are ”timely” attacks that take advantage of publicly available knowledge to craft contextually relevant emails. Using the example of spring break, attackers might send emails promising exclusive vacation deals or last-minute bookings at significantly reduced rates to lure recipients into clicking on malicious links. Another notable example observed by the cyber operations team is phishing emails disguised as job opportunities, primarily targeting students. These emails cunningly use external links or attached Word documents as bait. In a broader context, phishing emails regarding Duo MFA, credential renewals, or resets are crafted with urgency to deceive recipients across the organization. Another widespread tactic involves falsely alarming users about the de-provisioning of their email accounts if they fail to take immediate action. Faculty and staff are often the targets of these sophisticated phishing attempts. These include spoofed login pages from recognized domains (such as r20.rs6.net) and emails with forged usernames from external domains, employing lateral phishing strategies. To enhance their deceptive appearance, many phishing emails include images linked to the organization or Microsoft, exploiting the trust associated with these entities. This integration of social engineering techniques into phishing campaigns underscores the need for continual vigilance and education in cyber security practices.
III-C Security Incident Conclusion
The phishing email templates created for our comparative study are significantly influenced by the expert insights of our organization’s cyber operations team. In this research, we compare human and LLM phishing capabilities on three scenarios grounded in the team’s observations of past successful phishing campaigns. The first scenario represents a situation wherein a supervisor, leveraging the established rapport, emails their direct report, soliciting the execution of a specific task. This information is publicly available through organizational charts public directory and represents realistic information an attacker might use in lateral phishing. The second scenario unfolds through a more universally addressed email (e.g., events@ORG_NAME.edu), urging the recipient to undertake a particular action in relation to an impending event. The final scenario presents an interaction from an employee’s account, previously unassociated with the recipient, prompting an action aligned with a commonly observed phishing tactic within our organization (such as a fictitious job offer).
@@ III-C Security Incident Conclusion
  • Lateral phishing emails often mimic communications from a supervisor to their direct reports, presenting a deceptive appearance of legitimacy.
@@
  • The crafting of phishing emails may involve the use of LLMs.
III-D Research Questions
RQ1: How does the effectiveness of lateral phishing emails generated by Large Language Models compare to that of those crafted by humans within an educational organization? Understanding the effectiveness of lateral phishing attacks is critical to assessing organizational security vulnerabilities. This research question is novel as it investigates a large educational institution, contrasting with previous large-scale phishing research which predominantly focuses on commercial organizations. Our approach offers a unique perspective compared to the findings presented in previous prominent large scale phishing studies [14, 15]. Moreover, this question is crucial because it addresses the emerging threat of LLM-generated phishing emails, which can potentially lower the barrier for attackers by generating convincing, error-free content. By comparing the effectiveness of LLM and human-generated phishing emails, we can better understand the extent of this threat and inform more targeted defense strategies. Notably, our study compares LLM-generated emails against those crafted by a communications professional, representing the upper end of human ability in creating persuasive and coherent messages. This ensures a rigorous evaluation of the relative effectiveness of LLM-generated phishing emails against high-quality human-crafted emails.
RQ2: Which job roles and department affiliations exhibit the highest risk of falling victim to phishing emails, as evidenced by elevated click-through rates, and what patterns can be discerned to identify these high-risk groups? Identifying the job roles and department affiliations most susceptible to phishing attacks is essential for developing targeted awareness campaigns and allocating security resources effectively. By analyzing click-through rates across different job roles and departments, we can uncover patterns that may indicate specific vulnerabilities or knowledge gaps. This information is valuable for tailoring training programs and implementing more granular security measures to mitigate risks associated with high-risk groups. Furthermore, understanding these patterns can help organizations identify potential insider threats and develop strategies to address them proactively.
RQ3: What are the underlying reasons for individuals to engage with lateral phishing emails? Exploring the reasons and motivations behind individuals’ engagement with phishing emails is crucial for developing effective user education and awareness programs. By understanding the factors that influence users’ decision-making processes when faced with phishing attempts, organizations can design more targeted and persuasive training materials to help users identify and avoid these threats. Insights gathered from this investigation can inform the development of more comprehensive and adaptive defense strategies that account for the evolving nature of phishing attacks.
IV Threat Model